Suppose you have rights to gold mine for the next 20 years, during which you plan to extract 5000 kg every year. Assume a discount rate of 10%. The current price per kg is kshs.24, 000 but it’s expected to increase to 3% a year. Compute the present value of the gold.
1
Expert's answer
2019-04-03T09:04:11-0400

Present value of this annuity is:

"PV = \frac{P\times((1 - (1 + r)^{-n})}{r},"

"PV = \frac{24000\times5000\times((1 - (1 + 0.1)^{-20})}{0.1} = kshs1021627650."